# Towards a Science Exocortex

### v1

@@ 1 Introduction
While LLMs are often viewed purely as text-generators (e.g. for chat interactions), they have transformative potential owing to their ability to generate decisions and plans. For instance, LLMs can trigger software tools by providing them access to application programming interfaces (APIs).[42, 43, 44, 45, 46, 47, 48, 49, 50] Generations can be improved by inducing self-critique of output quality,[51, 52] or creating chains of thought through iterative self-prompting.[53, 54, 55] These systems can be turned into task-oriented autonomous agents by allowing them to iteratively propose and execute solutions.[46, 56, 57, 58, 59]
@@ 2.1 AI Agents
Research into LLM-based AI agents is ongoing, with several prototypes having been demonstrated.[46, 56, 57, 58, 61, 62] Although the optimal architecture remains an open question, current demos typically add several elements to the base LLM, such as: providing the LLM with access to various tools (software, web browsing, etc.), the ability to store information about its ongoing work (i.e. memory[63, 64, 65, 66]), some kind of loop to iterate on problems (inner monologue[67] or chain-of-thought[52, 53, 54, 55] or internal graph search[68]), and prompting suggesting breaking a problem into steps, and then working progressively on each step. Further improvements are possible using an architecture where task plans and status are captured in an explicit tree structure, which provides a flexible way to organize complex hierarchies.[59] Tree structures can be efficiently searched (e.g. Monte Carlo tree search) for reasoning and planning, yielding improvements in many tasks including math.[69, 70, 71, 72, 73]
@@
Additional research will be required to adapt agentic LLM approaches to scientific problems. Straightforward improvements would arise from training or fine-tuning LLMs on scientific documents, to ensure understanding of the relevant topics. Document retrieval can also easily improve LLM performance on scientific tasks.[37] Additional LLM specializations for science should also be considered. Golkar et al. proposed xVal, a specialized token encoding for numbers (scaling a dedicated embedding vector) which improves LLM handling of numerical tasks.[74] McLeish et al. used special positional embeddings (relative to start of number) and demonstrated vastly improved performance and generalization on simple arithmetic (addition and multiplication) tasks.[75] Xu et al. integrated symbolic expressions and logic rules into a chain-of-thought prompting strategy, demonstrating improved reasoning on logical tasks since the LLM was invoking formal logic and symbol manipulation during the solution.[55] Trinh et al. combined a language model with a symbolic solver to handle geometry theorems.[76] These kinds of approaches appear promising, suggesting that LLMs with slight adaptations could yield vastly improved reasoning for science and engineering tasks.
@@
An advantage of the exocortex architecture is that it can easily integrate more advanced AI agents as they are developed by others. In other words, we propose to separate the design/function of agents from their inter-communication, so that new agents can be added to the exocortex easily (by simply building a wrapper that supports the expected messaging between agents). The goal is to be able to leverage the growing ecosystem of AI modules being developed for science, including for simulating complex systems,[77] fluid dynamics,[78] crystallography,[79], chemistry tools,[80] protein representation[81] and design,[82] and pathology images.[83, 84]
@@ 2.1.4 Knowledge Mapping
A grand challenge in data science is to integrate data from disparate sources into a single model. Human scientists excel at this task, as they integrate insights provided from experimental data, calculations, literature they recall, and intuitions informed by years of scientific practice. When thinking about or discussing a complex topic, human scientists will naturally jump between different levels of abstraction and thus different scientific models. This combination of models (some highly quantitative, others heuristic) allows human scientists to compensate for the deficiencies in one model/reasoning by leveraging another. Approximating this efficient behavior in a synthetic knowledge system is challenging.
@@ 2.1.5 Literature Discovery
LLMs can also be exploited to automate tedious tasks. For instance, they can be used for ranking[200, 37] or classifying[37] scientific documents. This opens up a new possibility for researcher engagement with the literature, beyond the conventional activities of periodically searching for articles of interest and keeping a watch for relevant articles through networks (peers or automated). LLMs could be used to search, organize, rank, triage, and summarize papers, and thereby identify the most pertinent publications for human consideration.
@@ 2.2 Exocortex System
The complexity of interconnected agents, and the non-standardized (text-based) messaging between them poses a problem for automated monitoring, analysis, and optimization of these systems. On the other hand, it is possible that existing approaches for systems engineering can be recast to the context of AI swarms. For instance, machine-learning has benefiting enormously from gradient backpropagation,[239] which has essentially automated the process of optimizing complex neural network and AI models. By analogy, Yuksekgonul et al. proposed TextGrad as a text-based “differentiation” of AI systems.[240] Natural language feedback (e.g. criticism) of system outputs can be used as scores (analogous to loss), the variation in score as a function of changes in prompt can be used as a gradient, and gradients can be propagated across the system with knowledge of architecture. This allows automated optimization of LLM-interaction networks. Further developing techniques such as this may be crucial to properly optimzing exocortex-like systems.
@@ 2.2.1 AI-AI Interactions
LLMs generate ideas and decisions, but they are quite primitive in the sense that the ideas are reflexive rather than resulting from deep introspection.[238] The repeated waves of processing that occur within an LLM as it proceeds through tokens provides an opportunity to build-up more complex assessments, with the current understanding represented as updates to the residual stream. Improved behavior can thus be elicited by inducing the model to explicitly output reasoning steps.[241, 242] Interestingly, introducing even meaningless filler tokens into the output provides improved performance,[243] presumably owing to the additional computation cycles that are invoked. And yet, LLMs implement a relatively primitive and unidirectional method of thinking, as they are unable to revise the serialized output. Multiple research efforts aim to improve this by introducing a sort of deliberation cycle, such as by triggering self-critique of output,[51, 52] or generating chains of thought through iterative self-prompting.[53, 54, 55] Exploiting tree search (e.g. Monte Carlo) can further improve quality, especially on math problems.[69, 70, 71, 72] For scientific applications, versions of these methods that explicitly invoke formal logic are especially attractive.[55] One can also provide pre-designed thought-templates to improve reasoning on selected tasks;[244] building a catalog of templates for scientific tasks would be beneficial.
@@
The optimal architecture for providing LLMs with deliberative capabilities remains an open and exciting research question.[238] Current scaling suggests that LLMs have untapped potential that could be unlocked with appropriate designs. In parallel with algorithmic research, we propose that scientific researchers can make progress by simply expending compute to compensate for architectural weaknesses. For instance, consider tasking an agent-swarm with a problem, whereupon the agents generate ideas, ask each other questions, generate random permutations by combining ideas, rank all ideas, and only present the best ideas to the human. This workflow is highly wasteful in the sense that the majority of the generated content is never seen and indeed low-quality. Yet this invisible content can be viewed as the system’s internal deliberation. Even if this process is inefficient, its automated and unobtrusive nature can make the outputs sought-after by humans. In the context of science expenditures, the associated costs could be small relative to the value. There are tentative reports that this kind of extended search can yield substantial results.[258, 259, 260] We thus propose increasing investigation by physical scientists of such brute-force workflows, for generating content useful to researchers.

### v2

@@ 1 Introduction
While LLMs are often viewed purely as text-generators (e.g. for chat interactions), they have transformative potential owing to their ability to generate decisions and plans. For instance, LLMs can trigger software tools by providing them access to application programming interfaces (APIs).[42, 43, 44, 45, 46, 47, 48, 49, 50] Generations can be improved by inducing self-critique of output quality,[51, 52, 53] or creating chains of thought through iterative self-prompting.[54, 55, 56] These systems can be turned into task-oriented autonomous agents by allowing them to iteratively propose and execute solutions.[46, 57, 58, 59, 60]
@@ 2.1 AI Agents
Research into LLM-based AI agents is ongoing, with several prototypes having been demonstrated.[46, 57, 58, 59, 62, 63, 64, 65, 66] Although the optimal architecture remains an open question, current demos typically add several elements to the base LLM, such as: providing the LLM with access to various tools (software, web browsing, etc.), the ability to store information about its ongoing work (i.e. memory[67, 68, 69, 70]), some kind of loop to iterate on problems (inner monologue[71] or chain-of-thought[52, 54, 55, 56] or internal graph search[72]), and prompting suggesting breaking a problem into steps, and then working progressively on each step. Further improvements are possible using an architecture where task plans and status are captured in an explicit tree structure, which provides a flexible way to organize complex hierarchies.[60] Tree structures can be efficiently searched (e.g. Monte Carlo tree search) for reasoning and planning, yielding improvements in many tasks including math.[73, 74, 75, 76, 77]
@@
Additional research will be required to adapt agentic LLM approaches to scientific problems. Straightforward improvements would arise from training or fine-tuning LLMs on scientific documents, to ensure understanding of the relevant topics. Fine-tuning on math examples can elicit latent mathematical abilities.[78, 79] Document retrieval can also easily improve LLM performance on scientific tasks.[37] Additional LLM specializations for science should also be considered. Golkar et al. proposed xVal, a specialized token encoding for numbers (scaling a dedicated embedding vector) which improves LLM handling of numerical tasks.[80] McLeish et al. used special positional embeddings (relative to start of number) and demonstrated vastly improved performance and generalization on simple arithmetic (addition and multiplication) tasks.[81] Xu et al. integrated symbolic expressions and logic rules into a chain-of-thought prompting strategy, demonstrating improved reasoning on logical tasks since the LLM was invoking formal logic and symbol manipulation during the solution.[56] Trinh et al. combined a language model with a symbolic solver to handle geometry theorems.[82] Vashishtha et al. improved causal reasoning by providing axiomatic training examples.[83] These kinds of approaches appear promising, suggesting that LLMs with slight adaptations could yield vastly improved reasoning for science and engineering tasks.
@@
An advantage of the exocortex architecture is that it can easily integrate more advanced AI agents as they are developed by others. In other words, we propose to separate the design/function of agents from their inter-communication, so that new agents can be added to the exocortex easily (by simply building a wrapper that supports the expected messaging between agents). The goal is to be able to leverage the growing ecosystem of AI modules being developed for science, including for simulating complex systems,[84] optimizing differential equations,[85] fluid dynamics,[86] material discovery,[87] crystallography,[88], 2D materials,[89] chemistry tools,[90, 64] protein representation[91] and design,[92] and pathology images.[93, 94]
@@ 2.1.4 Knowledge Mapping
A grand challenge in data science is to integrate data from disparate sources into a single model. Human scientists excel at this task, as they integrate insights provided from experimental data, calculations, literature they recall, and intuitions informed by years of scientific practice. When thinking about or discussing a complex topic, human scientists will naturally jump between different levels of abstraction and thus different scientific models. This combination of models (some highly quantitative, others heuristic) allows human scientists to compensate for the deficiencies in one model/reasoning by leveraging another. Approximating this efficient behavior in a synthetic knowledge system is challenging. AI analogs of human synthesis would make knowledge integration explicit and documented, and provide integrated models for other software systems to leverage.
@@ 2.1.5 Literature Discovery
LLMs can also be exploited to automate tedious tasks. For instance, they can be used for ranking,[210, 37, 211] evaluating,[212] or classifying[37] scientific documents. This opens up a new possibility for researcher engagement with the literature, beyond the conventional activities of periodically searching for articles of interest and keeping a watch for relevant articles through networks (peers or automated). LLMs could be used to search, organize, rank, triage, and summarize papers, and thereby identify the most pertinent publications for human consideration.
@@
Another use for AI agents is to aid researchers in drafting scientific manuscripts. LLMs are fundamentally text-generation systems, and their role in productively generating long-form textual content is being extensively studied.[215, 216, 217] As often observed with LLMs, the quality of output can be improved through iteration, including using the LLM to generate an outline, self-critique output, and so on. There are additional challenges in using LLMs to generate scientific text, as consistency and correctness must not be compromised. Here too, there has been progress in using LLMs to automatically generate full-length technical documents.[218, 219, 220] The use of AI to generate text for inclusion in the scientific literature could be deleterious to science if the texts contain too many errors (compared to the human baseline). The exocortex design emphasizes the central role of the human researcher in assessing correctness and validating decisions/generations. We propose that the human researcher maintain the important role of validation, and thereby maintain responsibility for the quality of publications to which they attach their name.
@@ 2.2 Exocortex System
The complexity of interconnected agents, and the non-standardized (text-based) messaging between them poses a problem for automated monitoring, analysis, and optimization of these systems. On the other hand, it is possible that existing approaches for systems engineering can be recast to the context of AI swarms. For instance, machine-learning has benefiting enormously from gradient backpropagation,[257] which has essentially automated the process of optimizing complex neural network and AI models. By analogy, Yuksekgonul et al. proposed TextGrad as a text-based “differentiation” of AI systems.[258] Natural language feedback (e.g. criticism) of system outputs can be used as scores (analogous to loss), the variation in score as a function of changes in prompt can be used as a gradient, and gradients can be propagated across the system with knowledge of architecture. This allows automated optimization of LLM-interaction networks. Zhou et al. demonstrate how symbolic learning can be applied to optimizing LLM frameworks.[259] Further developing techniques such as these may be crucial to properly optimizing exocortex-like systems.
@@
The proposed exocortex architecture would treat agents as modules, allowing them to be swapped or for new agents to be added. Connecting agents to each other should enable progressively more complex automated workflows. However, a crucial open question is whether multi-agent workflows can scale to complex problems. For instance, even with a low per-step error rate, long task sequences could easily accumulate intolerable total error rates. The acceptable error-rate will be quite different for different parts of a workflow. For instance, imperfect ideas generated during ideation have low risk, as they will be identified and filtered out by the human easily (in human ideation there is value in initially considering erroneous ideas, as this can improve creativity). On the other hand, errors introduced by a data-analysis agent could be subtle and difficult to detect; yet errors in this stage would contaminate downstream analysis and thus invalidate the science. Errors in the experimental stage could waste valuable resources (time, experimental material, etc.) but are likely to be caught by human oversight. AI agent workflows can also be difficult to debug (owing to stochastic response) and brittle to maintain (changes in cloud models, changes in input data distribution, etc.). Thus, the scalability of multi-agent workflows is a crucial open question, requiring research and development. Frontier work in this area suggests that well-designed multi-step AI workflows may be able to generate coherent outputs.[219]
@@ 2.2.1 AI-AI Interactions
LLMs generate ideas and decisions, but they are quite primitive in the sense that the ideas are reflexive rather than resulting from deep introspection.[256] The repeated waves of processing that occur within an LLM as it proceeds through tokens provides an opportunity to build-up more complex assessments, with the current understanding represented as updates to the residual stream. Improved behavior can thus be elicited by inducing the model to explicitly output reasoning steps.[260, 261] Interestingly, introducing even meaningless filler tokens into the output provides improved performance,[262] presumably owing to the additional computation cycles that are invoked. And yet, LLMs implement a relatively primitive and unidirectional method of thinking, as they are unable to revise the serialized output. Multiple research efforts aim to improve this by introducing a sort of deliberation cycle, such as by triggering self-critique of output,[51, 52] or generating chains of thought through iterative self-prompting.[54, 55, 56] Exploiting tree search (e.g. Monte Carlo) can further improve quality, especially on math problems.[73, 74, 75, 76] For scientific applications, versions of these methods that explicitly invoke formal logic are especially attractive.[56, 83] One can also provide pre-designed thought-templates to improve reasoning on selected tasks;[263] building a catalog of templates for scientific tasks would be beneficial.
@@
The optimal architecture for providing LLMs with deliberative capabilities remains an open and exciting research question.[256] Current scaling suggests that LLMs have untapped potential that could be unlocked with appropriate designs. In parallel with algorithmic research, we propose that scientific researchers can make progress by simply expending compute to compensate for architectural weaknesses. For instance, consider tasking an agent-swarm with a problem, whereupon the agents generate ideas, ask each other questions, generate random permutations by combining ideas, rank all ideas, and only present the best ideas to the human. This workflow is highly wasteful in the sense that the majority of the generated content is never seen and indeed low-quality. Yet this invisible content can be viewed as the system’s internal deliberation. Even if this process is inefficient, its automated and unobtrusive nature can make the outputs sought-after by humans. In the context of science expenditures, the associated costs could be small relative to the value. There are tentative reports that this kind of extended search can yield substantial improvements.[277, 278, 279, 280, 281, 282, 283] We thus propose increasing investigation by physical scientists of such brute-force workflows, for generating content useful to researchers.
@@ 2.3.1 Automated Instruments
Conversely, it must be acknowledged that automation of scientific instruments and laboratory workflows represents a bottleneck for AI-driven science. While AI models and software can be rapidly iterated and improved, hardware system improvements are more capital-intensive and require longer-timescale design and construction efforts. Vendor-provided tools may use proprietary data formats and may not expose software interfaces that provide complete control of the system. These represent significant roadblocks to automation; the community should correspondingly demand commercial solutions that adhere to open data standards. Although mechanizing and automating laboratory work is by no means trivial, we argue that the value of any such effort will increase dramatically in the coming years, as AI agent control systems increase in sophistication.
@@ 3 Perspectives
The proposed multi-agent interactions and workflows highlight several open research questions. It is not known whether complex multi-step AI tasks will be sufficiently robust. The community must measure how AI capabilities scale, as a function of task complexity and inter-agent organizational architecture. The bottlenecks for scientific discovery—especially automated discovery workflows—must be elucidated. We speculate that LLMs will provide high utility for ideation and hypothesis generation, by providing the human with text digests and ranked ideas, and by acting as a conversational partner. However, integration with experimental tools is likely to lag, owing to the time and cost associated with building and testing laboratory automation systems. With respect to the overall exocortex system, we envision the largest roadblocks arising from managing the complexity of inter-communicating agents, and establishing sufficient reliability.